OpenStudy (anonymous):

The maximum speed of the Tupolev Tu-144 airliner is 694 m/s. What is the speed in kilometers per hour?

OpenStudy (trancenova):

m/s * 3600s/h * 1km/ m= 3600/1000m/s= 3.6 so m/s*3.6=km/h 694m/s= 3.6*694km/h=2498.4km/h